OpenStudy (anonymous):

1 over 4 + 4x over 5 = 11 over 20

OpenStudy (ranga):

Can you click on the button marked Equation below the reply box and type the question using the equation editor so the problem is clear?

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } + \frac{ 4x }{ 5 } =\frac{ 11 }{ 20 }\]

OpenStudy (anonymous):

\[\frac{ 4x }{ 5 }=\frac{ 11 }{ 20 }-\frac{ 1 }{4}\]

OpenStudy (anonymous):

\[\frac{ 4x }{ 5}=\frac{ 11-5 }{ 20 } = \frac{ 6 }{ 20 }\]

OpenStudy (anonymous):

x=\[x=3/8\]

OpenStudy (anonymous):

Thank u!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (ranga):

Click on "best Response" against mdasm. And close the question if it has been answered to your satisfaction.